### v2

# Cygnus X-3 as semi-hidden PeVatron

Abstract
Context. The high-mass X-ray binary Cygnus X-3 has been suggested for a long time to be a source of high-energy photons and neutrinos.
Aims. In view of the increased sensitivity of current experiments, we examine the acceleration and interactions of high-energy cosmic rays (CRs) in this binary system, assuming that the compact object is a black hole.
Methods. Using a test-particle approach in a Monte-Carlo framework, we employ as the basic CR acceleration mechanisms magnetic reconnection or 2.nd order Fermi acceleration and diffusive shock acceleration.
Results. We find that in all three scenarios CRs can be accelerated beyond PeV energies. High-energy photons and neutrinos are produced as secondaries in photo-hadronic interactions of CRs on X-ray photons and in the scattering on gas from the wind of the companion star. Normalising the predicted photon flux to the excess flux observed by LHAASO at energies above PeV in the direction of Cygnus X-3, a CR acceleration efficiency of $10^{-3}$ is sufficient to power the required CR luminosity. Our results suggest that the PeV photon flux from Cygnus X-3  could be in a bright phase significantly increased relative to the average flux of the last years.
Key Words.:
Cygnus X-3, high-mass X-ray binaries, multi-messenger astronomy
1 Introduction
Cygnus X-3, a high-mass X-ray binary in the direction of the Cygnus OB association, has had an outstanding impact on the development of gamma-ray astronomy. In 1983, Samorski & Stamm found a periodically modulated flux of neutral particles with energies above $2\times 10^{15}$ eV coincident with Cygnus X-3 in archival data from the Kiel air-shower array. This apparent detection and the supporting evidence provided during the following years by various experiments operating in the northern hemisphere triggered an enormous theoretical interest in this source (Gaisser & Stanev, 1985; Berezinsky et al., 1985, 1986). More importantly, these early results on Cygnus X-3 stimulated the building of air-shower arrays designed specifically for gamma-ray astronomy. However, none of these newly built arrays, like CASA-MIA and SPASE, detected a signal from Cygnus X-3, for a review of these early results see Protheroe (1994). Moreover, theoretical arguments suggest that most of these claims, in particular those by underground experiments, were erroneous as discussed already by Berezinsky et al. (1986). Still, it is a tantalising option that the initial claims for a photon signal were correct. In this case, Cygnus X-3 might have been in a bright phase in the early 1980s, which ended before the start of data-taking of the following generation of experiments. Such a strong variability is consistent with the decrease by at least a factor of 100 reported in the long-term average flux between the mid-1970s and mid-1980s (Bhat et al., 1986).
Around 40 years later, the advent of advanced air-shower arrays with improved hadron-photon separation power has started a new era in very-high energy gamma-ray astronomy. In particular, LHAASO has detected high-energy photons with energies up to few PeV in the direction of the Cygnus X superbubble (Cao et al., 2024). Moreover, the measured photon flux from the central region of the bubble is enhanced, with two out of eight PeV photons located inside a region of radius $0.5^{\circ}$. This excess, which corresponds to an increased flux by a factor of 5–10 compared to the average from the Cygnus superbubble, points towards additional point sources in this central region. The aim of this work is to examine if, and under which conditions, Cygnus X-3, which is located in the central region of the Cygnus superbubble, may be responsible for this excess PeV photon flux.
Using a phenomenological test-particle approach, we employ depending on the level of magnetisation in the jet different mechanisms for the acceleration of cosmic rays (CRs): In the strong magnetic turbulence close to the black hole (BH), we assume that 2.nd order Fermi acceleration or magnetic reconnection are efficient acceleration processes, while we suppose that diffusive shock acceleration (DSA) operates in the environment with small magnetisation at the termination shock of the two jets emanating from the BH. We perform simulations for two states of Cygnus X-3, which differ among others by the spectral energy distribution of X-ray photons. In the first one, called $S_{1}$, the jet has switched on right after the quiescent state, interacting with the surrounding medium close to the BH. In the second one, called $S_{2}$ and suggested by Koljonen et al. (2023) as a promising state for the production of high-energy secondaries, the jet has been on for some time, working against the Wolf-Rayet medium, forming a cocoon and a termination shock. Thus the soft X-ray state $S_{1}$ is most likely disk-dominated, while the hard X-ray state $S_{2}$ corresponds to a corona/jet-dominated state. We simulate the acceleration and interactions of CRs in a Monte-Carlo framework, including both hadronic and photo-hadronic interactions. For the high-energy photons which are produced as secondaries in these interaction we include absorption via pair production in the photon fields present.
The plan of this article is as follows: We start in Section 2 with a brief description of Cygnus X-3, summarizing our choice for the values of the relevant physical parameters which enter in the following calculations. Then we describe in Section 3 the acceleration scenarios we assume to operate in Cygnus X-3, and the Monte-Carlo scheme we employ to follow the time evolution of particles. Finally, we present in Section 4 the resulting time scales of the relevant acceleration, interactions and energy loss processes, and our numerical results for the fluxes of high-energy photons and neutrinos. We finish by a discussion of the results and the underlying assumptions, before we conclude.
2 Environment of Cyg X-3
Binary system
The determination of radial velocity curves for the Cygnus X-3 binary system is complicated by strong optical extinction and the difficult separation of wind features from those arising in the photosphere of the companion. Therefore, the individual masses of the components, $M_{\rm CO}$ of the compact object and $M_{\rm WR}$ of the companion Wolf-Rayet (WR) star, have remained uncertain up-to-date. Published results span a wide range from a BH with mass close to or above $20\,M_{\odot}$ (Schmutz et al., 1996; Hjalmarsdotter et al., 2008) down to upper limits of $3.6\,M_{\odot}$ (Stark & Saia, 2003). In contrast, the mass ratio of the binary stars can be constrained more precisely as $M_{\rm CO}/M_{\rm WR}=0.24\pm 0.06$ combining radial velocity curves derived from FeXXVI emission lines with infrared HeI absorption lines (Hanson et al., 2000). Luckily, several parameters important for the determination of the spectra of high-energy particles have a rather weak dependence on the masses of the binary system. For instance, the orbital separation $a$, which in turn influences the target density of stellar photons and the gas in the stellar wind, depends only as $a\propto(M_{\rm CO}+M_{\rm WR})^{1/3}$ on the total mass of the binary system. To be able to compare easier with earlier studies, we will use values close to the upper mass range discussed in the literature, setting $M_{\rm CO}=M_{\rm BH}=20M_{\odot}$ and $M_{\rm WR}=50M_{\odot}$ corresponding to the high-mass solution of  Szostek & Zdziarski (2008). We will disuses later how our results change if these masses are reduced. Thus the Schwarzschild radius of the BH is $R_{s}=6\times 10^{6}\,$cm, the orbital separation of the stars in the binary system is $a=3\times 10^{11}$ cm, while we use 9.6 kpc as distance to Cygnus X-3 (Reid & Miller-Jones, 2023).
Wind
The gas in the stellar wind of its WR companion provides an important target for the production of secondary particles in hadronic collisions. Mass conservation implies for the density profile of a spherically symmetric wind
$n(r)=\frac{\dot{M}_{\rm wind}}{4\pi r^{2}\mu m_{p}v_{\rm W}}$ (1)
where we use $\dot{M}_{\rm wind}=0.6\times 10^{-5}M_{\odot}$/yr and as wind velocity $v_{\rm W}=2\times 10^{8}$cm/s (Szostek & Zdziarski, 2008; Vilhu et al., 2021). For a proton rich environment, $\mu\simeq 1$, while for a helium rich environment, as suggested first by van Kerkwijk et al. (1992), it is $\mu\simeq 4$.
Photon fields
Another important target for the production of secondary particles are background photons from the stellar light of the WR star, from thermal emission of the accretion disk of the BH and from the synchrotron corona. These photon fields lead in photo-hadronic interactions to the production of secondary photons and neutrinos, and provide a target for the fragmentation of helium nuclei into nucleons via photo dissociation. For the thermal photons from the WR star, we use a Planck distribution with the temperature $T=10^{5}$ K which we rescale as $(R_{\rm WR}/r)^{2}$ at the distance $r$ to the WR star with radius $R_{\rm WR}=6\times 10^{10}$ cm (Sahakyan et al., 2014).
Figure 1: Spectral density of photon backgrounds in the state $S_{1}$ (left) and $S_{2}$ (right).
The accretion rate $\dot{M}$ in a X-ray binary may be strongly time-dependent, implying that also the nature of the accretion disk and thus the temperature profile of the disk varies. Since we are mainly interested in the bright phase of Cygnus X-3 with accretion close to the Eddington rate, we use a geometrically thin, optically thick Keplerian accretion disk (Shakura & Sunyaev, 1973; Chakrabarti, 1996). In this case, the thermal emission from the accretion disk is described by the temperature profile
$T(r)=\left(\frac{3GM\dot{M}}{8\sigma\pi r^{3}}\left[1-(R_{0}/r)^{1/2}\right]\right)^{1/4}.$ (2)
Since the radial extension of the disk is rather small, the total emission is close to a Planck distribution with $T(r)\simeq T(R_{0})$. For a helium-rich composition and the standard value $\eta=0.1$ for the accretion efficiency $\eta=L_{\rm Edd}/(\dot{M}_{\rm Edd}c^{2})$, the Eddington luminosity becomes $L_{\rm Edd}\simeq 2.6\times 10^{38}(M_{\rm BH}/M_{\odot})$erg/s which in turn fixes $\dot{M}=\lambda\dot{M}_{\rm Edd}$ for a chosen $\lambda$. The observed X-ray luminosity favours a large value of $\lambda$ and $M_{\rm BH}$ (Veledina et al., 2024). We calculate the density of disk photons at a given point in the jet integrating the photon emissivity into the corresponding solid angle ${\rm d}\Omega$ over the accretion disk (Cerutti et al., 2011), choosing $\lambda=0.5$ and as smallest radius $R_{0}$ of the acceleration and of the emission region six Schwarzschild radii, $R_{0}=6R_{s}\simeq 3.6\times 10^{7}\,$cm, while we set $R_{\rm out}=10R_{s}$ for the outer radius of the disk.
Finally, we have to fix the spectral density of X-ray photons in the corona. We identify the emission zone where energetic electrons are accelerated and radiate these photons via synchrotron and inverse Compton scattering with the acceleration region of hadrons. Then we use the INTEGRAL measurements from Cangemi et al. (2021) for the quiet state $S_{1}$ and for the state $S_{2}$ to determine the spectral number density of X-ray photons in the range 10-100 keV. In order to connect them to observations in the radio range from AMI-LA and RATAN (Piano et al., 2012), we assume an additional break at $E=1$ eV. For the size of the emission region, we assume $L=8R_{S}$ in the case of $S_{1}$ and $L=\tan\vartheta R_{S_{2}}$ with $R_{S_{2}}=7\times 10^{11}$ cm as the distance to the termination shock and $\vartheta=12^{\circ}$ as the jet opening angle (Mioduszewski et al., 2001; Spencer et al., 2022).
The resulting spectral number density of stellar, accretion disk and X-ray photons are shown in Fig. 1 for the state $S_{1}$ (left panel) and $S_{2}$ (right panel).
Magnetic field
We use as radial profile for the magnetic field strength around the BH
$B(r)=B_{0}\left(\frac{r_{0}}{r}\right)^{\delta},$ (3)
setting $B_{0}=2.1\times 10^{7}$ G at the jet injection point $r_{0}=10R_{s}\simeq 6\times 10^{7}$ cm (Miller-Jones et al., 2004). The slope $\delta$ is rather uncertain, with $0.5\leq\delta\leq 0.83$ being considered by Miller-Jones et al. Here, we follow Koljonen et al. (2018) who argued that $\delta\simeq 0.65$ fits best data.
In addition to the field strength, the ratio between the energy density in the magnetic field and the total energy density of the plasma is an important parameter. This ratio, the so-called magnetisation $\sigma_{\rm m}=B^{2}/(4\pi\rho)$ with $\rho\simeq n\gamma mc^{2}$ as the enthalpy density, becomes of order one at $r\simeq 10^{11}$ cm. Since the Alfvén velocity $v_{A}=c\sqrt{\sigma_{\rm m}/(1+\sigma_{\rm m})}$ approaches the speed of light for $\sigma_{\rm m}\gg 1$, the borderline between weak and strong magnetisation separates also the parameter space where DSA is more efficient than acceleration by second-order Fermi process or magnetic reconnection.
3 Theoretical framework
3.1 Acceleration scenarios
[…]
$t_{\rm sh}^{-1}=\frac{1}{E}\frac{{\rm d}E}{{\rm d}t}_{\rm sh}\simeq\zeta^{-1}\frac{v_{\rm sh}^{2}}{D}\simeq\frac{3Bv_{\rm sh}^{2}}{\zeta c{\cal R}}\simeq\eta\frac{cB}{{\cal R}}.$ (4)
With $\zeta\simeq 10-20$ for a parallel shock, $(v_{\rm sh}/c)^{2}\simeq 0.1$, we set conservatively $\eta=10^{-3}$. Note that we assume shock velocities $v_{\rm sh}$ small enough that relativistic effects like beaming or a large loss-cone of CRs at shock crossing can be neglected.
An alternative acceleration process which has recently attracted increased attention is magnetic reconnection. This acceleration mechanism has been applied to micro-quasars in general by de Gouveia Dal Pino & Lazarian (2005) and to Cygnus X-3 by Khiali et al. (2015). In this model, magnetic reconnection happens in current sheets produced when field lines arising from the accretion disk and of the BH magnetosphere encounter. If the reconnection velocity is fast enough, trapped particles between the two converging magnetic fluxes of opposite polarity can gain energy in a similar way to the first-order Fermi process, leading to (Kowal et al., 2012)
$t_{\rm rec}^{-1}=\frac{1}{E}\frac{{\rm d}E}{{\rm d}t}_{\rm rec}\simeq 1.3\times 10^{5}\left(\frac{{\cal R}}{m_{p}c}\right)^{-0.4}\frac{v_{A}}{L_{\rm rec}}$ (5)
with $v_{A}$ as the Alfvén velocity and $L_{\rm rec}$ as the extension of the reconnection zone. In this scenario, the acceleration region is close to the BH, at a distance between $\simeq 6R_{S}$ and $\simeq 10R_{S}$. As a result, the photon fields as potential targets for photo-hadronic interactions are much more intense than in the DSA case. Moreover, the magnetisation is very large and thus we assume that DSA is not effective.
Finally, we consider as alternative to magnetic reconnection 2.nd order Fermi acceleration in the strong magnetic turbulence close to the BH. For $\sigma_{\rm m}\gg 1$ and thus $v_{A}\simeq c$, the efficiency of 2.nd order Fermi acceleration is not suppressed relative to DSA. Moreover the slope of the produced particle distributions can become potentially universal, with $\alpha\sim 2.1$ found by  Comisso et al. (2024), and acceleration rate
$t_{\rm Fermi}^{-1}=\frac{1}{E}\frac{{\rm d}E}{{\rm d}t}_{\rm Fermi}\simeq 4\kappa(\Gamma\beta)^{2}\frac{v_{A}}{L_{c}}\simeq 10^{5}{\rm s}^{-1}.$ (6)
Here, $\kappa$ is an efficiency factor determined by Comisso & Sironi (2019) using PIC simulations as $\kappa\sim 0.1$, $|\boldsymbol{u}|=\Gamma|\boldsymbol{\beta}|\sim 1$ is the spatial part of the four-velocity of the scattering centers, and $L_{c}$ the coherence length of the turbulent magnetic field. We have assumed that the energy of the magnetic turbulence is mainly contained in large-scale modes, and have used $L_{c}=10^{6}$ cm for the numerical estimate. Thus these two processes lead for the chosen parameters roughly to the same maximal energy, cf. with the left panel of Fig. 2, and the same secondary production. However, second-order Fermi acceleration has the advantage that it can operate also at larger distances from the BH, as long as the jet is strongly magnetised. Since at larger distances both the acceleration efficiency and the secondary production would be reduced, the results from this acceleration mechanism would interpolate between those of magnetic reconnection very close to the BH and DSA at the termination shock. Thus we consider in our simulations only these two extreme cases.
Figure 2: Rates for synchrotron losses $t^{-1}_{\rm syn}$, photo-hadronic $t^{-1}_{p\gamma}$ and hadronic $t^{-1}_{pp}$ interactions of proton as function of energy compared to the acceleration rate by reconnection $t^{-1}_{\rm rec}$ and 2.nd order Fermi acceleration $t^{-1}_{\rm Fermi}$ in the state $S_{1}$ (left) and by DSA $t^{-1}_{\rm sh}$ in the state $S_{2}$ (right).
3.2 Monte Carlo scheme
We use a “leaky-box scheme” where for a chosen time step $\Delta t$ first the escape probability of a charged particle with energy $E$ from the acceleration region is calculated,
$p_{\rm esc}=1-\exp[(1-\alpha)\ln(1+\xi)].$ (7)
Here, $\alpha$ is the theoretically predicted slope of the differential energy spectrum of accelerated particles in the case of no interactions and energy losses, which we choose for both acceleration mechanisms as $\alpha=2.3$, while $\xi=({\rm d}E/{\rm d}t)_{\rm acc}\Delta t/E$ is the energy fraction gained according to Eqs. (4) and (5). If the particle stays in the acceleration process, the true energy gain is calculated as the sum of the acceleration gain and the continuous energy losses, $[({\rm d}E/{\rm d}t)_{\rm acc}-({\rm d}E/{\rm d}t)_{\rm CEL}]\Delta t$, where the latter include synchrotron and, for muons, inverse Compton losses. Finally, it is decided if the particle decays or interacts: First, the probability that something happens is compared to a random number $r$ uniformly distributed in [0:1[[0:1[[ 0 : 1 [, then a second random number $r$ is compared to the branching ratios for the relevant decay and interaction channels.
If a photo-hadronic interactions occurs, we use the modified version of the SOPHIA program (Mucke et al., 2000) developped by Kachelrieß et al. (2008) to generate secondary particles, while we employ QGSJET-IIc (Ostapchenko, 2011, 2013) for hadronic interactions. We take into account the finite decay time of all unstable particles except neutral pions and include only stable charged all particles in the acceleration process. For the decay of unstable particles, we employ SIBYLL (Fletcher et al., 1994).
After the particles escaped from the acceleration region, we assume that the magnetic field strength is small enough that they move ballistically. Except for photons, the interaction depth is small, $\tau\ll 1$, and multiple interactions can be neglected. Thus we add after escape a final interaction for hadrons with probability $p=1-\exp(-\tau)$ and let all unstable particles decay. In the case of photons, we calculate the pair-production probability on background photons in the direction of the observer. In addition, we add the pair-production probability on photons of the cosmic microwave background (CMB) and the extragalactic background light from Franceschini et al. (2008) to obtain the final flux observed on Earth, $F=F_{0}\exp(-\tau_{\gamma\gamma})$. We neglect the absorption on the starlight in the Milky Way, which adds only a minor correction relative to the absorption on the CMB (Vernetto & Lipari, 2016) and the effect of electromagnetic cascades what is well justified above the pair-production threshold, $E_{\gamma}\simeq(1-10)$ GeV.
Note that the acceleration time is much smaller than the orbital period of the system. Thus there is an orbital modulation of the fluxes, both because the interaction depth varies with the orbital phase and, in addition for photons, because the absorption probability changes. In the energy range we are mainly interested in, $(10^{14}$$10^{16})$ eV, the modulation is however only minor: The photon absorption probability is dominated by the CMB, while the interaction depths vary only by a factor a few, cf. Lammert (2025) for details. In the following, we present therefore results for the average fluxes.
4 Numerical results
We will start to discuss the case of magnetic reconnection and DSA assuming that protons are injected into the acceleration process, as it has been the common assumption in previous works. As the material close to the BH originates from the helium-rich wind of the WR companion star, we discuss after that briefly the case of helium as CR primaries.
Figure 3: Particle fluxes of CR protons, sum of all neutrino flavours and photon fluxes as function of energy $E$ for the case of reconnection in the state $S_{1}$ (left) and DSA in the state $S_{2}$ (right).
4.1 Proton injection
Having specified the acceleration mechanisms and the physical parameters determining interaction and energy loss processes, we can compare the resulting acceleration, interaction and energy loss rates. This will allow us to check and to interpret the numerical results from our Monte Carlo simulations. In Fig. 2, we show these rates for protons [^1] See Lammert (2025) for a discussion of other long-lived particles., comparing in the left panel the case of acceleration by reconnection in the state $S_{1}$ to the case of DSA in the state $S_{2}$ shown on the right. The dense photon fields close to the BH make photo-hadronic interactions to the most important energy-loss process, limiting the maximal energy of protons in the reconnection case to $\simeq 6\times 10^{16}$ eV. Hadronic interactions on gas from the stellar wind of the WR star play a negligible role during the acceleration phase, even if one would assume as Khiali et al. (2015) an increased gas density close to the BH. In the case of DSA in the state $S_{2}$, synchrotron losses and photo-hadronic interactions are of similar importance, with a slight excess of the former. The maximal energy of protons is somewhat smaller, $\simeq 3\times 10^{16}$ eV, but still high enough to expect secondaries with energies above PeV. Note also that in both cases the Hillas criterium allows higher CR rigidities than the interaction and energy losses: e.g., in $S_{1}$, the Hillas criterium allows for rigidites up to 300 PV.
We normalise the fluxes such that the CR luminosity satisfies $L_{\rm CR}=\eta_{\rm CR}\lambda L_{\rm Edd}\simeq 3.9\times 10^{38}$erg/s, i.e. we assume optimistically that 15% of the available energy is used to accelerate hadrons. In Fig. 3, we show the rescaled flux $E^{2}\Phi=E^{2}{\rm d}N/({\rm d}A{\rm d}t{\rm d}E)$ of CR protons (blue line), of the sum of all neutrino flavours (orange line) and of photons (red line) as function of energy $E$. In addition to the proton flux after interactions, the pressumed power-law $E^{-2.3}$ and the flux escaping the acceleration region are presented. The high-energy cutoffs agree with the expectations from Fig. 2, i.e. a rather soft suppression above $8\times 10^{16}$  eV in the state $S_{1}$ and a sharper suppression above $2\times 10^{16}$  eV in the state $S_{2}$. Similarly, the secondary fluxes in the state $S_{1}$ are as expected larger than in the state $S_{2}$. The red dotted line indicates the unabsorbed photon flux, while the dot-dashed line shows the photon flux taking into account pair production: At energies above $10^{14}$ eV pair production on CMB photons dominates, while at lower energies the absorption on stellar photons becomes important.
In Fig. 4, we show a close-up of the photon fluxes, where we split the flux into two components, depending on their production inside or outside the acceleration zone. The outer component is mainly produced in hadronic interactions, and thus the slope of this component repeats the slope of the proton flux at an energy a factor $\simeq 20$ higher. In addition, we show in Fig. 4 the gamma-ray flux measured by LHAASO in the direction of the Cygnus X superbubble from Cao et al. (2024). Note that Cygnus X-3 contributes only a fraction, mostly at the highest energies, of the total flux from the Cygnus superbubble. Since no detailed information on the exposure is published, we can only roughly estimate the integrated flux $\Phi(>E)$ from the central region above PeV energies as $\Phi(>{\rm PeV})=N/(AT)\simeq 9\times 10^{-14}/({\rm m^{2}s})$ using as effective area $A\simeq 10^{6}$m2 from Ma et al. (2022), $T\simeq 6000$ h for 3 years of data taking, and $N=2$ events above PeV. This corresponds to a photon flux of order $E^{2}\Phi(E={\rm PeV})\simeq 10^{-11}$eV/cm2s. Thus the obtained photon flux, which is in both states on the level $E^{2}\Phi(E={\rm PeV})\simeq 10^{-9}$eV/cm2s, should be scaled down by $10^{-2}$, what would reduce the required CR acceleration efficiency down to $\eta_{\rm CR}\sim 10^{-3}$. At this level, the photon flux would be well below the lower limits from the MAGIC collaboration (Aleksić et al., 2010). The correspondingly down-scaled neutrino flux is a factor $100$ below the 90% C.L. upper limit set by the IceCube collaboration (Abbasi et al., 2022). Even for $\eta_{\rm CR}=0.1$, the number of expected muon neutrino events/year in IceCube above 10 TeV is only around one in $S_{1}$ and two in $S_{2}$.
Figure 4: Photon fluxes as function of energy $E$ for the case of reconnection in the state $S_{1}$ (left) and by DSA in the state $S_{2}$ (right) compared to the gamma-ray flux measured by LHAASO (Cao et al., 2024) in the direction of the Cygnus X superbubble.
4.2 Helium injection
Next we discuss briefly the case of helium as CR primaries. Since the acceleration and diffusion of CRs depends only on rigidity, the rates given in Eqs. (4) and (5) apply directly to helium. As synchrotron losses for the same primary energy scale with $(q/m)^{4}$, the losses of helium are reduced by a factor 16. In the case of hadronic interactions, the He-He cross section is a factor 6–7 larger than the proton-proton cross section, while the number of targets is reduced by a factor of four for a fixed wind density $\rho$. Thus the secondary production due to hadronic interactions is increased by a factor 1.5–1.75 relative to the proton case discussed before. The differences are more pronounced in the case of interactions with background photons. Here, in addition to the usual photo-hadronic channel, the photo dissociation of helium nuclei into nucleons sets in at a $\simeq 10$ times lower threshold energy. Thus helium nuclei at the highest energies will fragment into nucleons, resulting into the same energy cutoff as obtained in the proton case. At lower energies, mainly helium nuclei will escape from the acceleration zone, leading to a somewhat increased secondary production. Since the interaction depth is small, the largest part of the accelerated helium nuclei escapes and contributes to the helium part of the Galactic CR spectrum.
5 Discussion
Let us now briefly review some of our main assumptions and discuss how our results change if we modify them.
  • Our values for the masses of the binary stars are at the higher end of the range discussed in the literature, and the choice $M_{\rm BH}=5M_{\odot}$ and $M_{\rm WR}=20M_{\odot}$ would be more in line with the analysis of, e.g., Koljonen & Maccarone (2017). This value for $M_{\rm BH}$ would raise the temperature of the accretion disk by only $\sim 30\%$, and we have seen that the disk photons play only a minor role as target. More importantly, the number density of X-ray photons in the state $S_{1}$ would increase by a factor $\simeq 4^{3}=64$, reducing the maximal energy in the $S_{1}$ state by a factor of order ten. Such a reduction of the maximal proton energy would mean that Cygnus X-3 is barely a PeV photon source in the $S_{1}$ state. Alternatively, a two-zone model would have to be considered where the X-ray emission is partly decoupled from the acceleration of hadrons.
  • The accretion rate $\dot{M}$ regulates the total amount of power available, of which the fraction $\eta_{\rm CR}\lambda$ can be channeled into the acceleration of CRs. X-ray observations favour super-Eddington accretion rates and/or a rather massive BH (Veledina et al., 2024). Moreover, radio observations show a very large variability in the radio flux (Gregory et al., 1972; Green et al., 2025). If this variability is connected to variations in $\dot{M}$ or $\eta_{\rm CR}$, then the PeV photon flux should be in an one-zone model correlated with the variation in the radio band. If, on the other hand, a smaller jet-opening angle and larger jet velocities are characteristic for flares (Spencer et al., 2022), then the PeV photon flux could be in the state $S_{2}$ anti-correlated to the radio flares.
  • In the state $S_{2}$, a combined scenario with both DSA at the recolliminaton or termination shock and 2.nd order Fermi throughout the jet or magnetic reconnection close the the BH could be possible. In particular, for a smaller value of $\delta$, the maximal energy achievable in DSA is reduced and this region would not contribute to high-energy photons. However, this region might be still the main contributor to the X-ray emission via leptonic processes. Hence such a case would be a realisation of a two-zone model, where the target density for photo-hadronic interactions is strongly reduced, and rather small BH masses would be unproblematic even for the acceleration via reconnection close to the BH.
  • We have negelected relativistic beaming effects. As the flux scales with the Doppler factor $\delta$ as $\delta^{3}$, already modest gamma factors could lead to a strong increase of the observed photon fluxes.
  • We have considered a purely hadronic model for the photon emission of Cygnus X-3, without trying to explain self-consistently the X-ray emission by electrons. A combined model where the acceleration of electrons and hadrons is simulated based on the same mechanism would help to constrain physical parameters like the magnetic field strength and the size of the acceleration region, but is deferred to a future work.
6 Conclusions
We have examined the acceleration and interactions of high-energy CRs in the high-mass X-ray binary Cygnus X-3, motivated by the recent observations of two PeV photons in the direction of Cygnus X-3. We have found that in all the three acceleration scenarios considered—magnetic reconnection. 2.nd order Fermi acceleration on magnetic turbulence and diffusive shock acceleration—CRs can be accelerated beyond PeV energies. In the hadronic scenario studied by us, high-energy photons are mainly produced in interactions with gas from the wind of the WR companion star. The CR luminosity to explain the excess flux observed by LHAASO at energies above PeV energies is rather low, requiring a CR acceleration efficiency of $10^{-3}$$10^{-2}$. This suggest that the PeV photon flux from Cygnus X-3  could be in a bright phase significantly increased relative to the average flux of the last years.
Acknowledgements.
We would like to thank Karri Koljonen for useful discussions and helpful comments on the draft, and Egor Podlesnyi for help in estimating the expected number of IceCube events. M.K. is grateful to the late Venya Berezinsky for introducing him to the story of Cygnus X-3 in the 1980s.